# Arithmetical Complexity and Absoluteness of Rigidity Phenomena for Ulam Sequences

### v1

@@ 6.1 Absoluteness for arithmetical sentences
Lemma 6.1 (Absoluteness for arithmetical sentences).
Let $M$ and $N$ be transitive models of $\mathrm{ZFC}$ with $\mathbb{N}^{M}=\mathbb{N}^{N}=\mathbb{N}$. Then for every arithmetical sentence $\varphi$ in the language of arithmetic we have
This is exactly Lemma 2.3 from Section 2, and its proof rests on the fact that the satisfaction relation for arithmetical formulas is primitive recursive in the underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$); see, for instance, [Shoenfield1967, Chapter VI] or [Kunen1980, Chapter VIII].
@@ 6.3 Reverse-mathematical remarks
Thus the arguments of Theorems 4.2, 4.4, 5.3, and 5.8 can be formalised in $\mathsf{RCA}_{0}$ without change, yielding the stated equivalences. ∎
@@ 7.1 Model-theoretic background
A basic example is Presburger arithmetic: the complete theory of $(\mathbb{N},+,0,1)$ is NIP and dp-minimal, and its definable sets are semilinear.
@@ 7.3 Rigidity and tameness
We now record some immediate model-theoretic consequences of Lemma 7.6. Recall that definitional expansions do not change the class of definable sets (up to naming additional definable predicates) or the interpretability power of a theory: any structure interpretable in a definitional expansion is already interpretable in the original structure.
Proposition 7.7 (Sample model-theoretic consequence).
@@
Assume that $\mathrm{Rig}(a,b)$ holds and that the associated pattern code $c$ witnesses rigidity in the sense of Lemma 7.6. Then the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is a definitional expansion of Presburger arithmetic. In particular:
@@
By Lemma 7.6, under the assumption of $\mathrm{Rig}(a,b)$ there exists a formula $\varphi(x)$ in the language $\{0,1,+\}$ with parameters such that for all $m\in\mathbb{N}$,
@@
Thus the predicate symbol $\mathrm{U}_{a,b}$ is definable in the reduct $(\mathbb{N},+,0,1)$, and $\mathcal{U}_{a,b}$ is a definitional expansion of $(\mathbb{N},+,0,1)$: every formula in the expanded language can be translated into an equivalent formula in the original Presburger language by replacing atomic instances of $\mathrm{U}_{a,b}(x)$ with $\varphi(x)$.
Let $T_{\mathrm{Pres}}$ denote the complete theory of $(\mathbb{N},+,0,1)$ and let $T_{U}$ denote $\mathrm{Th}(\mathcal{U}_{a,b})$. The previous paragraph shows that $T_{U}$ is a conservative definitional extension of $T_{\mathrm{Pres}}$. Therefore any structure interpretable in $T_{U}$ is already interpretable in $T_{\mathrm{Pres}}$.
It is a classical fact that Presburger arithmetic does not interpret the full ring $(\mathbb{N},+,\times)$; in particular, multiplication is not interpretable in $(\mathbb{N},+,0,1)$ (see e.g. discussions of Presburger arithmetic in standard model-theory texts). If $(\mathbb{N},+,\times)$ were interpretable in $T_{U}$, then by definitionality it would be interpretable in $T_{\mathrm{Pres}}$, a contradiction. This proves that $T_{U}$ does not interpret $(\mathbb{N},+,\times)$.
Finally, it is well-known that Presburger arithmetic is NIP and dp-minimal: its definable sets are semilinear and admit a simple combinatorial classification. Since NIP and dp-minimality are preserved under definitional expansions (any independence pattern in the expanded language would translate to one in the original language), it follows that $\mathcal{U}_{a,b}$ is again NIP and dp-minimal. In particular, every family of definable subsets of $\mathbb{N}$ in $\mathcal{U}_{a,b}$ has uniformly bounded VC-dimension.
In particular, since Presburger arithmetic is decidable, so is $\mathrm{Th}(\mathcal{U}_{a,b})$ under the hypothesis of $\mathrm{Rig}(a,b)$: any first-order sentence in the expanded language can be effectively translated into an equivalent Presburger sentence and decided using standard decision procedures (e.g. via Cooper’s algorithm or automata-theoretic methods). ∎
@@
There is a substantial body of work on decidable and undecidable expansions of Presburger arithmetic; see, for example, Point’s study of decidable extensions [Point2000DecidablePresburger] and B‘es’ survey on arithmetical definability [Bes2001SurveyArithDef].
7.4 Questions and conjectures
We conclude with a few natural questions suggested by the preceding discussion. We state them informally, as precise formulations would require additional background.
  • NIP under weaker rigidity. The definability result in Lemma 7.6 requires the full strength of $\mathrm{Rig}(a,b)$ as formulated in Section 4. Is it already sufficient, for NIP (or for non-interpretability of multiplication), to assume weaker forms of rigidity, such as eventual periodicity of the gap sequence together with mild boundedness conditions on the error terms?
  • Potential wild behaviour. Suppose there exists a pair $(a,b)$ for which $U(a,b)$ fails to be regular and does not satisfy any strong rigidity statement of the above type. Can the corresponding expansion $\mathcal{U}_{a,b}$ exhibit model- theoretic wildness? For instance:
    • Does there exist a coprime pair $(a,b)$ such that the expansion $\mathcal{U}_{a,b}=(\mathbb{N},+,0,1,\mathrm{U}_{a,b})$ has the independence property?
    • More modestly, can $\mathcal{U}_{a,b}$ interpret the full ring $(\mathbb{N},+,\times)$ or at least a nontrivial initial segment of true arithmetic, *without* any strong rigidity hypothesis on $U(a,b)$?
    Conversely, do natural combinatorial constraints on Ulam sequences prevent such behaviour? Outside the rigidity context, very little is known about the model-theoretic behaviour of $\mathcal{U}_{a,b}$. Our arithmetisation results show that the predicate $\mathrm{U}_{a,b}$ is recursive and that $(\mathbb{N},+,\mathrm{U}_{a,b})$ has a decidable elementary diagram, but this alone does not preclude highly wild classification-theoretic behaviour.
  • Uniform families. The family rigidity statement $\mathrm{FRig}$ suggests studying the theory of the multi-sorted structure whose sorts are $(\mathbb{N},+)$ and a parameter space for $(a,b)$, together with a uniform predicate for $\mathrm{U}(a,b,m)$. Under $\mathrm{FRig}$, can one obtain a uniform model- theoretic classification of the family $(\mathcal{U}_{a,b})_{(a,b)}$, e.g. a uniform NIP bound or a uniform VC-dimension bound for certain definable families?
  • Connections with known expansions of Presburger arithmetic. Expansions of Presburger arithmetic by unary predicates (for example, subsets of $\mathbb{N}$ with prescribed asymptotic structure) have been studied in various contexts. To what extent can existing classification results for such expansions be applied to $\mathcal{U}_{a,b}$, given the combinatorial information we obtain about $U(a,b)$?
These questions lie beyond the scope of the present paper, but they indicate that Ulam sequences and sets may provide a rich testing ground for the interaction between additive combinatorics and the model theory of expansions of Presburger arithmetic.

### v2

@@ 6.1 Absoluteness for arithmetical sentences
This is exactly Lemma 2.3 from Section 2, and its proof rests on the fact that the satisfaction relation for arithmetical formulas is primitive recursive in the underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$); see, for instance, [Shoenfield1967, Chapter VI]; for a related discussion of definability and absoluteness in set theory, see [Kunen1980, Chapter IV and its appendices].
@@ 6.3 Reverse-mathematical remarks
Thus the arguments of Theorems 4.3, 4.6, 5.3, and 5.8 can be formalised in $\mathsf{RCA}_{0}$ without change, yielding the stated equivalences. ∎
@@ 7.1 Model-theoretic background
A basic example is Presburger arithmetic: the complete theory $\mathrm{Th}(\mathbb{N},+,0,1)$ of addition on the natural numbers is NIP and dp-minimal, and its definable sets are semilinear.
@@ 7.3 Rigidity and tameness
We now develop the model-theoretic consequences of various structural conditions on Ulam sequences. We begin by introducing a hierarchy of combinatorial assumptions, ranging from strong rigidity to weaker density conditions.
Definition 7.7 (Hierarchy of structural conditions).
Let $(a,b)$ be coprime with $a<b$. We consider the following conditions on $U(a,b)$, listed roughly from strongest to weakest:
  1. (R1)
    Strong rigidity: There exist $c,N$ such that for all $m\geq N$, $m\in U(a,b)$ iff $\mathrm{InPattern}(c,a,b,m)$.
  2. (R2)
    Gap regularity: The gap sequence $(g_{k})$ is eventually periodic, i.e. $\exists N\,\exists p\,\forall k\geq N\,(g_{k+p}=g_{k})$.
  3. (R3)
    Bounded gaps: There exists $B$ such that $g_{k}\leq B$ for all $k$.
  4. (R4)
    Density existence: The limit $d(U(a,b))=\lim_{n\to\infty}|U(a,b)\cap[0,n]|/n$ exists.
  5. (R5)
    Positive lower density: $\underline{d}(U(a,b))=\liminf_{n\to\infty}|U(a,b)\cap[0,n]|/n>0$.
The implications among these conditions are as follows:
$\text{(R1)}\Rightarrow\text{(R2)}\Rightarrow\text{(R3)}\Rightarrow\text{(R5)},\qquad\text{(R2)}\Rightarrow\text{(R4)}.$
Verification of implications.
[…]
$|U(a,b)\cap[0,n]|\geq\frac{n-a}{B}+1,$
giving $\underline{d}(U(a,b))\geq 1/B>0$. ∎
Remark 7.8.
Note that (R3) does not imply (R4) in general: bounded gaps ensure positive lower density but do not by themselves guarantee that the limit $\lim_{n\to\infty}|U(a,b)\cap[0,n]|/n$ exists. Similarly, (R4) does not imply (R5): there exist infinite sets with asymptotic density zero (e.g. the squares or the primes). It is not known whether any Ulam sequence has density zero; all computed examples suggest positive density, but a proof that $d(U(a,b))>0$ whenever it exists remains open.
We now establish the key consequences of gap regularity.
Proposition 7.9 (Consequences of gap regularity).
Suppose $U(a,b)$ satisfies condition (R2): the gap sequence $(g_{k})$ is eventually periodic with period $p$ and threshold $N$. Then:
  1. (i)
    The density $d(U(a,b))$ exists and equals
    $d(U(a,b))=\frac{p}{\sum_{j=0}^{p-1}g_{N+j}}.$
  2. (ii)
    The set $U(a,b)$ is the union of a finite set and a set that is Presburger-definable with parameters.
  3. (iii)
    The expansion $\mathcal{U}_{a,b}$ is NIP.
Proof.
(i) For $k\geq N$, write $k=N+mp+r$ with $0\leq r<p$ and $m\geq 0$. The partial sums of gaps satisfy
$u_{N+mp+r}=u_{N}+m\cdot G+\sum_{j=0}^{r-1}g_{N+j},$
where $G=\sum_{j=0}^{p-1}g_{N+j}$ is the sum of one period of gaps. For large $n$, the number of Ulam elements in $u_{N},n]$ is asymptotically $p\cdot(n-u_{N})/G$. It follows that
$\lim_{n\to\infty}\frac{|U(a,b)\cap[0,n]|}{n}=\frac{p}{G}.$
(ii) The tail $\{u_{k}:k\geq N\}$ consists of elements of the form
$u_{N}+mG+\sum_{j=0}^{r-1}g_{N+j}$
for $m\geq 0$ and $0\leq r<p$. For each fixed $r$, this is an arithmetic progression with first term $u_{N}+\sum_{j=0}^{r-1}g_{N+j}$ and common difference $G$. Thus the tail is a union of $p$ arithmetic progressions. Arithmetic progressions are Presburger-definable, so the tail is Presburger-definable with parameters $(u_{N},g_{N},\dots,g_{N+p-1})$. The finite initial segment $\{u_{k}:k<N\}$ is also Presburger-definable (as a finite set of constants).
(iii) By (ii), there is a formula $\psi(x,\bar{c})$ in the language $\{0,1,+\}$ with a finite tuple of parameters $\bar{c}$ such that $m\in U(a,b)$ iff $(\mathbb{N},+,0,1)\models\psi(m,\bar{c})$. Naming the parameters $\bar{c}$ gives a (definitional) expansion of Presburger arithmetic that remains NIP (since NIP is preserved under naming constants). Now adding a new unary predicate symbol $P(x)$ together with the defining axiom $\forall x\,(P(x)\leftrightarrow\psi(x,\bar{c}))$ is again a definitional expansion, so the resulting theory is NIP. Interpreting $P$ as $\mathrm{U}_{a,b}$ yields $\mathcal{U}_{a,b}$, which is therefore NIP. ∎
Remark 7.10 (Gap regularity vs. strong rigidity).
Lemma 7.6 shows that strong rigidity (R1) implies Presburger-definability of $U(a,b)$ (with parameters), and hence the tameness of $\mathcal{U}_{a,b}$ discussed below. Part (ii) of Proposition 7.9 shows that one actually needs only gap regularity (R2) to obtain Presburger-definability of the tail and therefore of $U(a,b)$ up to a finite set. In particular, the hypothesis of (R2) in Proposition 7.11 below is strictly weaker than the strong rigidity assumption used in Lemma 7.6.
The extra content of strong rigidity is effectivity: it provides an explicit pattern code $c$ from which the Presburger definition of $U(a,b)$ can be computed, whereas gap regularity merely asserts the existence of an eventual periodicity pattern without specifying it in terms of interval-with-mask decompositions. In principle, (R2) is strictly weaker than (R1): one could imagine a sequence with eventually periodic gaps but where the positions of the Ulam elements within each period vary in a way not captured by a simple pattern code. In practice, all known gap-regular examples also satisfy strong rigidity, and it remains open whether (R2) can hold without (R1).
@@
We now record the main model-theoretic consequence. Recall that definitional expansions do not change the class of definable sets (up to naming additional definable predicates) or the interpretability power of a theory: any structure interpretable in a definitional expansion is already interpretable in the original structure.
Proposition 7.11 (Model-theoretic tameness under gap regularity).
@@
Assume that $U(a,b)$ satisfies gap regularity (R2) of Definition 7.7, i.e., its gap sequence $(g_{k})$ is eventually periodic. Then the theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is a definitional expansion of Presburger arithmetic. In particular:
@@
By Proposition 7.9(ii), under the assumption of gap regularity (R2) there is a formula $\varphi(x)$ in the language $\{0,1,+\}$ with parameters such that for all $m\in\mathbb{N}$,
@@
In other words, $U(a,b)$ is definable (up to a finite initial segment, which is also definable) in the underlying Presburger structure. Thus the predicate symbol $\mathrm{U}_{a,b}$ is definable in the reduct $(\mathbb{N},+,0,1)$, and $\mathcal{U}_{a,b}$ is a definitional expansion of $(\mathbb{N},+,0,1)$: every formula in the expanded language can be translated into an equivalent formula in the original Presburger language by replacing atomic instances of $\mathrm{U}_{a,b}(x)$ with $\varphi(x)$.
Let $T_{\mathrm{Pres}}$ denote the complete theory of $(\mathbb{N},+,0,1)$ and let $T_{U}$ denote $\mathrm{Th}(\mathcal{U}_{a,b})$. Since $T_{U}$ is a conservative definitional extension of $T_{\mathrm{Pres}}$, any structure interpretable in $T_{U}$ is already interpretable in $T_{\mathrm{Pres}}$.
It is a classical fact that Presburger arithmetic does not interpret the full ring $(\mathbb{N},+,\times)$; in particular, multiplication is not interpretable in $(\mathbb{N},+,0,1)$. If $(\mathbb{N},+,\times)$ were interpretable in $T_{U}$, then by definitionality it would be interpretable in $T_{\mathrm{Pres}}$, a contradiction. This proves that $T_{U}$ does not interpret $(\mathbb{N},+,\times)$.
Finally, Presburger arithmetic is NIP and dp-minimal, and NIP and dp-minimality are preserved under definitional expansions. Hence $\mathcal{U}_{a,b}$ is again NIP and dp-minimal, and, in particular, every family of definable subsets of $\mathbb{N}$ in $\mathcal{U}_{a,b}$ has uniformly bounded VC-dimension. Since Presburger arithmetic is decidable, the same holds for $\mathrm{Th}(\mathcal{U}_{a,b})$. ∎
In Section 7.4 we ask how far one can push such tameness without assuming strong rigidity or even gap regularity.
@@
There is a substantial body of work on decidable and undecidable expansions of Presburger arithmetic; see, for example, Point’s study of decidable extensions [Point2000DecidablePresburger] and Bès’ survey on arithmetical definability [Bes2001SurveyArithDef].
7.4 Unconditional and partial results
The preceding results establish model-theoretic tameness of $\mathcal{U}_{a,b}$ under the hypothesis of gap regularity (R2), and in particular under strong rigidity (R1). In this subsection we investigate what can be said without these assumptions, either unconditionally or under weaker structural hypotheses. While full tameness results remain out of reach in the general case, we identify several constraints that the Ulam construction imposes on the expansion $\mathcal{U}_{a,b}$ and formulate precise conjectures about intermediate cases.
7.4.1 The unique-sum property and its consequences
The defining feature of Ulam sequences—that each term beyond the initial pair has a unique representation as a sum of two distinct earlier terms—is a strong structural constraint. We begin by formalising this and extracting some unconditional consequences.
Definition 7.13 (Representation function).
For a set $P\subseteq\mathbb{N}$ and $n\in\mathbb{N}$, the representation function $r_{P}(n)$ counts the number of ways to write $n$ as a sum of two distinct elements of $P$:
$r_{P}(n)=\bigl|\{(x,y)\in P^{2}:x<y\text{ and }x+y=n\}\bigr|.$
Lemma 7.14 (Representation bound for Ulam sequences).
Let $(a,b)$ be coprime with $a<b$. For every $m\in U(a,b)\setminus\{a,b\}$ we have $r_{U(a,b)}(m)=1$. Moreover, for every $n\notin U(a,b)$ with $n>b$, either $r_{U(a,b)\cap[1,n)}(n)=0$ or $r_{U(a,b)\cap[1,n)}(n)\geq 2$.
Proof.
The first claim is immediate from the definition of the Ulam sequence: an integer $m>b$ is adjoined to $U(a,b)$ at stage $k$ if and only if $m$ is the smallest integer exceeding all previously included terms and having exactly one representation as a sum of two distinct earlier terms. Thus $r_{U(a,b)\cap[1,m)}(m)=1$ at the moment of inclusion, and since no further terms less than $m$ are added subsequently, this remains $r_{U(a,b)}(m)=1$ in the final sequence.
For the second claim, suppose $n>b$ and $n\notin U(a,b)$. At the stage when the Ulam construction considers $n$, either $n$ has no representations as a sum of two distinct earlier terms (so $r_{U(a,b)\cap[1,n)}(n)=0$), or $n$ has at least two such representations (so $r_{U(a,b)\cap[1,n)}(n)\geq 2$). In the latter case, since the Ulam sequence $U(a,b)\cap[1,n)$ is already determined by stage $n$, this representation count persists. ∎
The representation bound suggests heuristic constraints on the density of Ulam sequences, though rigorous upper bounds appear difficult to establish.
Remark 7.15 (Density heuristics).
Intuitively, one expects the density of any Ulam sequence to be bounded well away from $1$. The reason is as follows: if $U=U(a,b)$ had density $d$ close to $1$, then the number of pairs from $U\cap[1,n]$ would be $\binom{dn}{2}\approx d^{2}n^{2}/2$, while the range of possible sums is only $[a+b,2n]$, an interval of length <2n<2n. By the pigeonhole principle, many integers would have multiple representations, and by Lemma 7.14, such integers cannot belong to $U$ (beyond the initial terms). This suggests a tension between high density and the unique-representation requirement.
Making this precise is nontrivial. A naive application of the Cauchy–Schwarz inequality to the second moment $\sum_{m}r_{U_{n}}(m)^{2}$ yields only weak bounds on the number of multiply-represented integers, insufficient to establish a sharp density constraint such as $\overline{d}(U)\leq 1/2$. The difficulty lies in controlling the distribution of representation counts among non-members: while each must have 0 or $\geq 2$ representations, the total “energy” absorbed by non-members depends on how representation counts are distributed, and crude bounds lose too much information.
Numerical evidence suggests that densities of Ulam sequences, when they exist, are quite small. For the classical sequence $U(1,2)$, computations indicate $d(U(1,2))\approx 0.07$ [Steinerberger2017HiddenSignal], and other well-studied pairs yield similarly small values. The gap between these empirical densities and even a conjectural bound of $1/2$ suggests that the rigidity phenomena impose much stronger constraints than naive counting arguments capture.
7.4.2 The additive character of the Ulam construction
We now turn to structural features of the expansion $\mathcal{U}_{a,b}$ that follow from the purely additive nature of the Ulam construction.
Definition 7.16 (Additively closed construction).
A subset $P\subseteq\mathbb{N}$ is said to arise from an additively closed construction if there is a recursive procedure defining $P$ in which membership of $n$ in $P$ depends only on:
  • the values of $P\cap[1,n)$ (the initial segment);
  • linear combinations and comparisons of elements of $P\cap[1,n)$;
  • bounded quantification over $P\cap[1,n)$.
In particular, the procedure never references multiplication, division, exponentiation, or other non-additive operations on the elements themselves.
Lemma 7.17.
For every coprime $(a,b)$, the Ulam sequence $U(a,b)$ arises from an additively closed construction in the sense of Definition 7.16.
Proof.
The Ulam construction proceeds as follows: given $U(a,b)\cap[1,n)$, an integer $n$ is adjoined if and only if (i) $n>\max(U(a,b)\cap[1,n))$, and (ii) there is exactly one pair $(x,y)$ with $\in U(a,b)\cap[1,n), x<yx<y$, and $x+y=n$. Condition (i) involves only comparison. Condition (ii) involves only addition of elements and counting (bounded quantification). No multiplicative operations appear. Thus the construction is additively closed. ∎
The significance of this observation is that additively closed constructions produce predicates whose interaction with the additive structure of $\mathbb{N}$ is constrained. While we cannot prove unconditionally that $\mathcal{U}_{a,b}$ fails to interpret multiplication, we can articulate why such an interpretation would be surprising.
Proposition 7.18 (Decidability of the atomic diagram).
Let $(a,b)$ be coprime with $a<b$. The atomic diagram of $\mathcal{U}_{a,b}=(\mathbb{N},+,0,1,\mathrm{U}_{a,b})$ is decidable: given any quantifier-free sentence $\varphi$ with parameters from $\mathbb{N}$, one can effectively determine whether $\mathcal{U}_{a,b}\models\varphi$.
Proof.
The predicate $\mathrm{U}(a,b,m)$ is $\Delta^{0}_{1}$ by Theorem 3.2, hence decidable. Any quantifier-free formula in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ with parameters involves only finitely many membership queries to $U(a,b)$ and finitely many linear comparisons, all of which are decidable. ∎
Remark 7.19 (On Presburger definability).
Presburger-definable subsets of $\mathbb{N}$ are exactly the eventually periodic sets: finite unions of arithmetic progressions and singletons. It is expected, in line with computational evidence and the conjectured lack of eventual periodicity of gaps for most pairs $(a,b)$, that $\mathrm{U}_{a,b}$ is not definable in $(\mathbb{N},+,0,1)$ without parameters unless strong rigidity holds. Indeed, definability of $U(a,b)$ in Presburger arithmetic would force eventual periodicity of its characteristic function, which appears to fail for generic pairs based on extensive numerical data.
Remark 7.20 (On quantifier elimination).
It seems unlikely that $\mathcal{U}_{a,b}$ admits quantifier elimination in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ when $U(a,b)$ is not eventually periodic. Already the set $\{(x,y)\in\mathbb{N}^{2}:x+y\in U(a,b)\}$ is definable by a simple existential formula, but describing it quantifier-free would require the predicate $\mathrm{U}_{a,b}$ to interact with addition in a highly structured way. We do not pursue a definitive result on quantifier elimination here.
7.4.3 Comparison with known interpretations of arithmetic
To place the model-theoretic status of $\mathcal{U}_{a,b}$ in context, we briefly recall when an expansion $(\mathbb{N},+,P)$ interprets full arithmetic.
Definition 7.21 (Multiplicative encoding).
We say that a predicate $P\subseteq\mathbb{N}$ encodes multiplication if there exist first-order formulas $\varphi_{\times}(x,y,z)$ and $\varphi_{D}(x)$ in the language $\{0,1,+,P\}$ such that $\varphi_{D}$ defines an infinite subset $D\subseteq\mathbb{N}$ and
$(\mathbb{N},+,P)\models\varphi_{\times}(a,b,c)\quad\Longleftrightarrow\quad a\cdot b=c$
for all $a,b,c\in D$. Equivalently, $(\mathbb{N},+,P)$ interprets a nontrivial fragment of $(\mathbb{N},+,\times)$.
Predicates defined in terms of strong multiplicative or exponential growth are known to produce undecidable expansions where arithmetic is interpretable. Classical examples include:
  • $P=\{k^{n}:n\in\mathbb{N}\}$ for a fixed base $k\geq 2$ (powers);
  • $P=\{n!:n\in\mathbb{N}\}$ (factorials);
  • $P=\{p_{n}:n\in\mathbb{N}\}$ (the sequence of primes).
For these and related results, see Villemaire [Villemaire1992], Muchnik–Semenov [MuchnikSemenov2003], Point [Point2000DecidablePresburger], and the survey by Bès [Bes2001SurveyArithDef]. The common feature of these examples is that the predicate $P$ encodes multiplicative information through its growth pattern or divisibility structure.
In contrast, “additively natural” predicates tend to yield tamer expansions.
Remark 7.22 (Sparse sets and tameness; cf. [MuchnikSemenov2003, Bes2001SurveyArithDef]).
Let $P\subseteq\mathbb{N}$ be a set of density zero satisfying suitable sparseness conditions (for example, $|P\cap[0,n]|=O(\log n)$). Under certain additional hypotheses on the growth rate and structure of $P$, the expansion $(\mathbb{N},+,P)$ remains decidable and does not interpret $(\mathbb{N},+,\times)$. The precise conditions vary across different results in the literature; we refer to [MuchnikSemenov2003, Bes2001SurveyArithDef] for details.
Ulam sequences occupy an intermediate regime: they have positive density (conjecturally), so the extreme sparseness results do not directly apply, but they lack the multiplicative structure of the examples above. This motivates the following conjecture.
Conjecture 7.23 (Non-interpretability of multiplication).
For every coprime $(a,b)$, the expansion $\mathcal{U}_{a,b}$ does not interpret $(\mathbb{N},+,\times)$. That is, there is no formula $\varphi(x,y,z)$ in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ defining the graph of multiplication on an infinite domain.
Evidence for this conjecture includes:
  • The purely additive character of the Ulam construction (Lemma 7.17);
  • The absence of any apparent mechanism by which the unique-sum property could encode multiplicative relations;
  • The contrast with known multiplication-encoding predicates, all of which involve exponential, factorial, or similar growth patterns foreign to Ulam sequences.
7.4.4 Weaker structural conditions
We now consider what can be said under conditions weaker than gap regularity (R2). Recall from Definition 7.7 that bounded gaps (R3), density existence (R4), and positive lower density (R5) form progressively weaker assumptions.
For the bounded-gaps case (R3), we expect tameness properties to hold but do not have complete proofs.
Conjecture 7.24 (Tameness under bounded gaps).
Suppose $U(a,b)$ satisfies condition (R3): there exists $B$ such that all gaps $g_{k}$ are at most $B$. Then:
  1. (i)
    The expansion $\mathcal{U}_{a,b}$ is NIP.
  2. (ii)
    For any formula $\varphi(x;y)$ in the language $\{0,1,+,\mathrm{U}_{a,b}\}$ with $|x|=1$, the VC-dimension of the family $\{\varphi(\mathbb{N};b):b\in\mathbb{N}^{|y|}\}$ is finite.
Remark 7.25.
Intuitively, bounded gaps should force some form of tameness: the predicate $\mathrm{U}_{a,b}$ stays within bounded distance of a union of arithmetic progressions (in the sense that between any two consecutive Ulam elements there are at most B1B-1 non-Ulam elements), and one expects the resulting definable families to have bounded VC-dimension. Results of Aschenbrenner–Dolich–Haskell–Macpherson–Starchenko [Aschenbrenner2016VapnikChervonenkis] on VC-density in NIP theories suggest that such bounds should exist, but verifying the hypotheses of their theorems for $\mathcal{U}_{a,b}$ under (R3) alone requires additional work that we do not undertake here.
We summarise the known and conjectured implications in the following table.
Condition NIP dp-minimal No $(\mathbb{N},+,\times)$
Strong rigidity (R1) Yes (Prop. 7.11) Yes (Prop. 7.11) Yes (Prop. 7.11)
Gap regularity (R2) Yes (Props. 7.97.11) Yes (Prop. 7.11) Yes (Prop. 7.11)
Bounded gaps (R3) Conjectured Unknown Conjectured
Density exists (R4) Unknown Unknown Conjectured
Positive lower density (R5) Unknown Unknown Conjectured
No assumption Conjectured Unknown Conjectured
7.4.5 A dichotomy conjecture
The preceding analysis suggests a possible dichotomy for expansions of Presburger arithmetic by Ulam predicates.
Conjecture 7.26 (Ulam dichotomy).
For every coprime $(a,b)$, exactly one of the following holds:
  1. (a)
    The expansion $\mathcal{U}_{a,b}$ is a reduct of a definitional expansion of Presburger arithmetic (possibly with countably many parameters), and in particular is NIP, dp-minimal, and does not interpret $(\mathbb{N},+,\times)$.
  2. (b)
    The expansion $\mathcal{U}_{a,b}$ interprets $(\mathbb{N},+,\times)$, and in particular has the independence property and is model-theoretically wild.
Moreover, case (b) does not occur: every Ulam expansion falls into case (a).
This conjecture asserts that the purely additive nature of the Ulam construction is incompatible with the encoding of multiplication, regardless of whether combinatorial rigidity holds. A proof would likely require new techniques combining additive combinatorics with definability theory.
Remark 7.27 (Relation to Zilber’s trichotomy).
Conjecture 7.26 is reminiscent of trichotomy and dichotomy principles in geometric model theory. In Zilber’s trichotomy for strongly minimal sets [Zilber1993], a strongly minimal set is either “trivial,” “group-like,” or “field-like,” with the latter corresponding to interpretability of algebraically closed fields. Our conjecture posits an analogous dichotomy for Ulam expansions: either the expansion remains “additively tame” or it becomes “arithmetically wild.” The content of the conjecture is that the second alternative never occurs for Ulam predicates.
7.4.6 Effectivity and computability considerations
We conclude this subsection with remarks on the effective content of the results under partial rigidity. The following is essentially a repackaging of the gap-regularity analysis in a more computational form.
Proposition 7.28 (Effective bounds under gap regularity).
Suppose $U(a,b)$ satisfies gap regularity (R2) with computable witnesses, i.e. there is an algorithm that, given $(a,b)$, outputs a threshold $N$ and period $p$ such that $g_{k+p}=g_{k}$ for all $k\geq N$. Then:
  1. (i)
    The density $d(U(a,b))$ is computable (as a rational number) from the witnesses.
  2. (ii)
    The theory $\mathrm{Th}(\mathcal{U}_{a,b})$ is decidable.
  3. (iii)
    Membership in any definable subset of $\mathbb{N}^{k}$ is decidable, uniformly in the defining formula.
Proof.
(i) From $N$ and $p$, compute $G=\sum_{j=0}^{p-1}g_{N+j}$ by simulating the Ulam construction up to stage $N+p$. Then $d(U(a,b))=p/G$ by Proposition 7.9(i).
(ii) By Proposition 7.9(ii), the predicate $\mathrm{U}_{a,b}$ is Presburger-definable with parameters determined by the witnesses. A formula in the language of $\mathcal{U}_{a,b}$ can therefore be translated into an equivalent Presburger formula (with the parameter values substituted). Presburger arithmetic is decidable (e.g. via Cooper’s algorithm or automata-theoretic methods), so the translated formula can be decided.
(iii) Given a formula $\varphi(\bar{x})$ in the language of $\mathcal{U}_{a,b}$ and a tuple $\bar{n}\in\mathbb{N}^{k}$, we can decide $\mathcal{U}_{a,b}\models\varphi(\bar{n})$ by the procedure in (ii). ∎
Remark 7.29.
The hypothesis of “computable witnesses” in Proposition 7.28 is stronger than mere truth of (R2). It is conceivable that $U(a,b)$ is gap-regular but that no algorithm can find the threshold $N$ and period $p$. In that case, the density and theory would still be determined (by absoluteness), but not necessarily effectively computable. The rigidity programme of Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019JNT] aims in part to provide such computable witnesses for specific families.